OpenStudy (anonymous):

Please help!!! (1/3-1/9) to the 2nd power

OpenStudy (anonymous):

its 26/81 but try simplifying

OpenStudy (anonymous):

the answer is 4/81 but how did u come up with that

OpenStudy (anonymous):

how do u work the problem out?

OpenStudy (anonymous):

i didnt i used a fraction calculator

OpenStudy (anonymous):

oh I need to know how to work the problem out

OpenStudy (anonymous):

ohh ok you take 1/3 - 1/9^2 so 9 is the only one that has the 2 power

OpenStudy (anonymous):

then what do u do

OpenStudy (anonymous):

1-1 = 1 9*9 = 81 3- 81 = 78 so its 1/78

OpenStudy (anonymous):

its saying the answer is 4/81 but how did they get that answer

OpenStudy (anonymous):

that was in the answer key in the back of the book but it doesnt show how it is done

OpenStudy (anonymous):

\[\left(\begin{matrix}1/3-1/9 \\ ?\end{matrix}\right)\]

OpenStudy (anonymous):

to the second power

OpenStudy (anonymous):

(3/9 - 1/9)^2 = (2/9)^2 = 4/81

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome